i just wanted to know how to calculate change in no. of moles of any reaction.


for example,


2C6H6(l)+15O2(g)---->12CO2(g)+6H2O(l)

Grade:11

2 Answers

Aman Bansal
13 years ago

dearv ivek,

AS IN YOUR EXAMPLE
NO. OF REACTANTS=2+15=17
NO. OF PRODUCTS=12+6=18

THEREFORE CHANGE IN NO. OF MOLES=01

BEST OF LUCK

We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL

Vivek Balakrishnan
13 years ago

actually the answer is -3/2. please let me know about this. And i generally want to know how to calculate the no. of moles in any type of reaction.